A solution prepared by bubbling chlorine gas through a diluted sodium hydroxide solution is described by the following equation:

            Cl2(g) + 2 OH-(aq) → Cl-(aq) + OCl-(aq) + H2O(l)

The chlorine needed for this reaction is obtained from the electrolysis of molten sodium chloride. How much volume of 0.30 M hypochlorite solution can you prepare from chlorine obtained from the electrolysis using 3.0 amp current for 25 minutes assuming 100% yield?
Answer #1

Mass obtained in electrolysis = Z×I×T

Z= molar mass/96500

I= current = 3A

T= Time = 25mins = 1500 seconds

Mass obtained = (71×3×1500)/96500

Mass of Chlorine = 3.31g

Moles of Chlorine = 3.31/71 = 0.0466

Moles of Chlorine = moles of hypochlorite

Concentration = moles /volume

Volume = moles /concentration

Volume = 0.0466/0.3M = 0.155L = 155mL

Volume required = 155mL